# HumanSplat: Generalizable Single-Image Human Gaussian Splatting with Structure Priors

### v1

@@ Abstract
Despite recent advancements in high-fidelity human reconstruction techniques, the requirements for densely captured images or time-consuming per-instance optimization significantly hinder their applications in broader scenarios. To tackle these issues, we present HumanSplat that predicts the 3D Gaussian Splatting properties of any human from a single input image in a generalizable manner. In particular, HumanSplat comprises a 2D multi-view diffusion model and a latent reconstruction transformer with human structure priors that adeptly integrate geometric priors and semantic features within a unified framework. A hierarchical loss that incorporates human semantic information is further designed to achieve high-fidelity texture modeling and better constrain the estimated multiple views. Comprehensive experiments on standard benchmarks and in-the-wild images demonstrate that HumanSplat surpasses existing state-of-the-art methods in achieving photorealistic novel-view synthesis. Project page: https://humansplat.github.io/
@@ 1 Introduction
Realistic 3D human reconstruction is a fundamental task in computer vision with widespread applications in numerous fields, including social media, gaming, e-commerce, telepresence, etc. Previous works for single-image human reconstruction can be broadly categorized into explicit and implicit approaches. Explicit methods pymaf2021 ; PIXIE:2021 ; alldieckTex2ShapeDetailedFull2019 ; alldieck2018dv , such as those based on parametric body models like SMPL SMPL:2015 ; SMPL-X:2019 , estimate human meshes by directly optimizing parameters and clothing offset to fit the observed image. However, these methods often struggle with complex clothing styles and require lengthy optimization. Implicit methods saito2019pifu ; saito2020pifuhd ; alldieck2022photorealistic ; hu2023sherf ; huang2022elicit represent humans using continuous functions, such as occupancy saito2019pifu , SDF park2019deepsdf , and NeRF mildenhall2021nerf . They are good at modeling flexible topology but are limited in scalability and efficiency due to the high computational cost of training and inference. Recent advances in 3D Gaussian Splatting kerbl20233d (3DGS) have provided a balance between efficiency and rendering quality for reconstructing detailed 3D human models, which however rely on multi-view zielonka2023drivable ; moreau2023human ; qian20233dgs or monocular video li2024gaussianbody ; hu2023gauhuman ; hu2023gaussianavatar inputs. Recent popular human reconstruction studies TECH:2023 ; HumanSGD:2023 ; humannorm:2023 ; HAVEFUN:2023 focus on the score distillation sampling (SDS) technique SDS:2022 to lift 2D diffusion priors to 3D, but time-consuming optimization (e.g., 2 hours) is required for each instance. Some generalizable and large-reconstruction-model-based works hu2023sherf ; ActorsNeRF:2023 ; HumanLRM:2024 ; GPS:2023 can directly generalize the regression of 3D representations but either disregard human priors or require multi-view inputs, limiting the stability and feasibility in downstream applications.
@@
In this paper, we propose a novel generalizable approach HumanSplat for single-image human reconstruction by introducing a generalizable Gaussian Splatting framework with a pre-trained 2D video diffusion model and well-designed 3D human structure priors. Different from the existing human 3DGS methods, our approach directly infers Gaussian properties from a single image, obviating the necessity for per-instance optimization or densely captured images and empowering to generalize effectively across diverse scenes while efficiently delivering high-quality reconstructions.
@@
The key insight behind our method is to reconstruct Gaussian properties from the diffusion latent space in a generalizable end-to-end architecture, integrating a 2D generative diffusion model as appearance prior and a human parametric model as structure prior. Specifically, facing this under-constraint reconstruction problem from single-view input with extensive invisible regions, we first leverage a 2D video diffusion model (denoted as novel-view synthesizer) to hallucinate the unseen parts of clothed humans. Then a generalizable latent reconstruction transformer is proposed to enable interaction among the generated diffusion latent embeddings and the geometric information of the structured human model, enhancing the quality of 3DGS reconstruction. During interactions using inaccurate human prior like the SMPL model, a projection strategy is designed to balance the robustness and flexibility, where we compensate for inaccuracies by projecting 3D tokens into 2D space and conducting searches within adjacent windows using project-aware attention. Another specific challenge of human reconstruction is to capture fine detail in visually sensitive areas, such as the face and hand. To address this problem, we exploit the semantic cues from the structure priors and propose semantics-guided objectives to further promote the fine details reconstruction quality. By amalgamating these crucial components into a unified framework, our method achieves state-of-the-art performance in striking the right balance between quality and efficiency.
@@
Figure 1: Our method achieves state-of-the-art rendering quality while maintaining the fastest runtime. (a) Qualitative results: LGM LGM:2024 and GTA GTA:2024 are generalizable but in lower quality, TeCH TECH:2023 exhibits issues with multi-face rendering and is time-consuming. In contrast, our method achieves higher fidelity in a much shorter time. (b) Performance and runtime comparison: metrics are evaluated on the challenging Twindom dataset.
@@ 2 Related Work
Single-Image Human Reconstruction. Single-image human reconstruction methods fall into explicit and implicit approaches. Explicit methods use parametric body models SMPL:2015 ; SMPL-X:2019 to estimate minimally clothed human meshes pymaf2021 ; PIXIE:2021 ; Kanazawa2018_hmr ; pare ; pymafx2023 ; li2023niki ; chen2022learning ; VIBE ; Kolotouros2019_spin , adding clothing via 3D offsets alldieckTex2ShapeDetailedFull2019 ; alldieck2018dv ; alldieck2018videoavatar ; alldieck2019peopleInClothing ; zhu2019hierarchMeshDeform ; lazova2019textures360 or garment templates bhatnagar2019multiGarmentNet ; jiang2020bcnet ; feng2023learning . However, these methods are constrained by topology, particularly with loose clothing. Implicit methods saito2019pifu ; saito2020pifuhd ; alldieck2022photorealistic ; han2023high ; albahar2023humansgd ; hu2023sherf ; huang2022elicit ; HumanLRM:2024 ; Zhang_2024_CVPR , utilizing representations like occupancy, signed distance fields (SDF) and Neural Radiance Fields (NeRF), offer flexibility with topology, allowing for accurate depiction of 3D clothed humans. Recent methods he2020geoPifu ; corona2023s3f ; xiu2022icon ; zheng2021pamir ; huangARCHAnimatableReconstruction2020 ; heARCHAnimationReadyClothed2021 ; cao2022jiff ; liao2023car ; yang2023dif ; cao2023sesdf ; zhang2023globalcorrelated ; robustfusion ; su2022robustfusionPlus ; xiu2022econ ; Jiang_2023_CVPR ; zheng2024ohta combine implicit representation with explicit human priors for better robustness. Among these, GTA zhang2023globalcorrelated uses transformers with fixed embeddings for translating image features into 3D tri-plane features. TeCH huang2024tech employs diffusion-based models for visualizing unseen areas but needs extensive optimization and accurate SMPL-X models. Another approach, HumanSGD albahar2023humansgd , uses diffusion models for texture inpainting but struggles with mesh inaccuracies from saito2020pifuhd . NeRF-based methods like SHERF hu2023sherf generate human NeRFs from single images using hierarchical features, while ELICIT huang2022elicit leverages CLIP CLIP:2021 for contextual understanding. These NeRF-based methods produce high-quality images but often struggle with detailed 3D mesh generation and require extensive optimization time.
@@
Human Gaussian Splatting. While implicit representations like SDF or NeRF struggle with balance optimization efficiency and rendering quality, the high efficiency of 3DGS kerbl3Dgaussians has advanced 3D human creation. Recent 3DGS methods have targeted multi-view videos, monocular video sequences, and sparse-view input. Multi-view video methods like D3GA zielonka2023drivable , HuGS moreau2023human , and 3DGS-Avatar qian20233dgs exploit dense spatial and temporal information for detailed models. Additionally, HiFi4G jiang2023hifi4g combine 3DGS with a dual-graph mechanism to preserve spatial-temporal consistency, ASH pang2023ash utilizes mesh UV parameterization for real-time rendering, and Animatable Gaussians li2023animatable improve garment dynamics via pose projection mechanism and 2D CNNs. The monocular video only provides sufficient observation as humans move, so the monocular human Gaussian Splatting methods li2024gaussianbody ; hu2023gauhuman ; hu2023gaussianavatar ; li2023human101 ; GVA ; SplattingAvatar ; GoMAvatar ; lei2023gart ; HAHA often resort to using LBS for mapping to a canonical space and introducing additional regularization terms. For sparse-view images, GPS-Gaussian GPS:2023 achieves high rendering speeds without per-subject optimization by using a feed-forward way with efficient 2D CNNs encoding from diverse 3D human scan data. In comparison, our work stands out as the first one-shot generalizable human GS method, using only single image input and achieving high-quality reconstruction without any optimization and fine-tuning.
@@
Generalizable Large Reconstrucion Model. Large reconstruction model (LRM) hong2023lrm ; tochilkin2024triposr presents that with sufficient parameters and training datasets, a deterministic feed-forward Transformer vaswani2017attention is capable of learning a triplane feature for volumetric rendering from a single-view image. Targeting general 3D object, TriplaneGaussian zou2023triplane combines 3DGS with triplane as a hybrid 3D representation for fast rendering, where point clouds are first extracted from an image and then projected on triplane features to decode 3DGS attributes. Most recent methods further leverage 2D diffusion models that can generate multi-view images simultaneously shi2023mvdream ; wang2023imagedream ; shi2023zero123++ ; li2023instant3d to provide plausible and abundant inputs for the following reconstruction, so they focus on the sparse multi-view reconstruction task. Instant3D li2023instant3d follows the technique of LRM and produces triplane features from 4 posed images by a Transformer encoder. CRM wang2024crm , MeshLRM wei2024meshlrm and InstantMesh xu2024instantmesh replaces triplane NeRF by FlexiCubes shen2023flexible to extract meshes directly. LGM tang2024lgm , GRM xu2024grm and GS-LRM zhang2024gslrm predict 3DGS attributes from each input pixel szymanowicz2024splatter ; charatan2024pixelsplat and combine the Gaussians from multiple viewpoints as the final 3D output. For human task, CharacterGen peng2024charactergen follows Instant3D li2023instant3d to adopt a two-stage approach: first, generating a 2D multi-view canonical images, then using multi-view SDF reconstruction for cartoon digital characters, without direct interaction between stages. A more compact and elegant method HumanLRM HumanLRM:2024 replaces training data of LRM hong2023lrm from general objects with human data to predict triplane NeRF, showcasing the capacities of the large model. While it relies heavily on the dataset, often leading to issues like missing limbs due to a lack of prior knowledge. In contrast, our 3DGS-based method leverages the latent diffusion and reconstruction model in an end-to-end framework and introduces human priors by addressing their inaccuracy with specific projection strategies, achieving better generalization and stability with even less data.
@@ 3 Method
Figure 2: Overview of HumanSplat. (a) Multi-view latent features are first generated by a fine-tuned video diffusion model (Novel View Synthesizer in Sec. 3.3). (b) Then, the Latent Reconstruction Transformer (Sec. 3.4) interacts global latent features (Sec. 3.4.1) and human geometric prior (Sec. 3.4.2). (c) Finally, the semantic-guided objectives (Sec. 3.5) are proposed to reconstruct the final human 3DGS.
@@ 3.1 Preliminary
SMPL SMPL:2015 is a widely-used parametric human body model that is created by skinning and blend shapes, and learned from thousands of 3D body scans. The SMPL model SMPL:2015 utilizes shape parameters $\boldsymbol{\beta}\in\mathbb{R}^{10}$ and pose parameters $\boldsymbol{\theta}\in\mathbb{R}^{24\times 3}$ to parameterize the deformation of the human body mesh $\mathcal{M}(\boldsymbol{\beta},\boldsymbol{\theta}):\boldsymbol{\beta}\times\boldsymbol{\theta}\mapsto\mathbb{R}^{6890\times 3}$.
3D Gaussians Splatting kerbl20233d is an recently proposed 3D representation that represents 3D content by a set of $\mathbf{N_{p}}$ colored Gaussians $\mathcal{G}=\{\mathbf{g}_{i}\}_{i=1}^{\mathbf{N_{p}}}$. Each Gaussian $\mathbf{g}_{i}=\exp\left(-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu}_{i})^{\top}\boldsymbol{\Sigma}_{i}^{-1}(\mathbf{x}-\boldsymbol{\mu}_{i})\right)$ has opacity $\boldsymbol{\sigma}_{i}\in\mathbb{R}$ and color $\mathbf{c}_{i}\in\mathbb{R}^{3}$ attributes, where $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$ is its location and $\boldsymbol{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$ implies its scaling $\mathbf{s}\in\mathbb{R}^{3}$ and orientation quaternion $\mathbf{q}\in\mathbb{R}^{4}$ in 3D space. These attributes together $\mathbf{G}\in\mathbb{R}^{\mathbf{N_{p}}\times 14}$ can represent a 3D object, and by projecting to the image plane with opacity-based color composition in the depth order, it can be differentiably rendered into 2D images in real-time.
@@ 3.2 Overview
Given a single input image of a human body $\mathbf{I_{0}}\in\mathbb{R}^{H\times W\times 3}$, our task is to reconstruct a 3D representation, i.e., 3DGS in this work, from which novel-view images can be rendered. As illustrated in Fig. 2, our method leverages a 2D diffusion model and a novel latent reconstruction transformer that adeptly integrates 2D appearance priors, human geometric priors, and semantic cues within a universal framework. First, we use SMPL estimator feng2021collaborative ; Pavlakos2019_smplifyx to predict human structure priors $\mathcal{M}$, and CLIP CLIP:2021 to generate the image embedding of the input image $\mathbf{c}$. A latent temporal-spatial diffusion model voleti2024sv3d , referred to as the novel-view synthesizer (Sec. 3.3), is fine-tuned and produces multi-view latent features $\left\{\mathbf{F}_{i}\in\mathbb{R}^{h\times w\times c}\right\}_{i=1}^{\mathbf{N}}$, where $\mathbf{N}$ is the number of views, and $h$, $w$ and $c$ are the height, width and channel of the latent features. Then, we utilize a novel latent reconstruction transformer (Sec. 3.4) that adeptly integrates human geometric prior and latent features within a universal framework, which predicts Gaussian attributes $\mathbf{\mathbf{G}}\coloneqq\left\{(\boldsymbol{\mu}_{i},\boldsymbol{q}_{i},\boldsymbol{s}_{i},\boldsymbol{c}_{i},\boldsymbol{\sigma}_{i})|i=1,...,\mathbf{N_{p}}\right\}$ and rendered into new images, where $\mathbf{N_{p}}$ represents the number of Gaussian points. Furthermore, to achieve high-fidelity texture modeling and better constrain the estimated multiple views, we designed a hierarchical loss that incorporates human semantic priors (See 3.5). Note that during training, the network is trained using 3D scan data to ensure accurate supervision from various viewpoints, where the registered SMPL is fitted via the multi-view version of SMPLify Pavlakos2019_smplifyx . During inference, the model employs PIXIE feng2021collaborative to predict human structure priors $\mathcal{M}$ and synthesize novel views from only the single image input based on the trained model, without any fine-tuning and optimization.

### v2

@@ Abstract
Despite recent advancements in high-fidelity human reconstruction techniques, the requirements for densely captured images or time-consuming per-instance optimization significantly hinder their applications in broader scenarios. To tackle these issues, we present HumanSplat, which predicts the 3D Gaussian Splatting properties of any human from a single input image in a generalizable manner. Specifically, HumanSplat comprises a 2D multi-view diffusion model and a latent reconstruction Transformer with human structure priors that adeptly integrate geometric priors and semantic features within a unified framework. A hierarchical loss that incorporates human semantic information is devised to achieve high-fidelity texture modeling and impose stronger constraints on the estimated multiple views. Comprehensive experiments on standard benchmarks and in-the-wild images demonstrate that HumanSplat surpasses existing state-of-the-art methods in achieving photorealistic novel-view synthesis. Project page: https://humansplat.github.io.
@@ 1 Introduction
Realistic 3D human reconstruction is a fundamental task in computer vision with widespread applications in numerous fields, including social media, gaming, e-commerce, telepresence, etc. Previous works for single-image human reconstruction can be broadly categorized into explicit and implicit approaches. Explicit methods [1, 2, 3, 4], such as those based on parametric body models like SMPL [5, 6], estimate human meshes by directly optimizing parameters and clothing offset to fit the observed image. However, these methods often struggle with complex clothing styles and require lengthy optimization. Implicit methods [7, 8, 9, 10, 11] represent humans using continuous functions, such as occupancy [7], SDF [12], and NeRF [13]. While they excel at modeling flexible topology, these methods are limited in scalability and efficiency due to the high computational cost associated with training and inference. Recent advances in 3D Gaussian Splatting (3DGS) [14] have provided a balance between efficiency and rendering quality for reconstructing detailed 3D human models, which however rely on multi-view image [15, 16, 17] or monocular video [18, 19, 20] input. Recent popular human reconstruction studies [21, 22, 23, 24] focus on the score distillation sampling (SDS) technique [25] to lift 2D diffusion priors to 3D, but time-consuming optimization (e.g., 2 hours) is required for each instance. Some generalizable and large-reconstruction-model-based works [10, 26, 27, 28] can directly generalize the regression of 3D representations but either disregard human priors or require multi-view inputs, limiting the stability and feasibility in downstream applications.
@@
In this paper, we propose a novel generalizable approach HumanSplat for single-image human reconstruction by introducing a generalizable Gaussian Splatting framework with a fine-tuned 2D multi-view diffusion model and well-conceived 3D human structure priors. Different from existing human 3DGS methods, our approach directly infers Gaussian properties from a single input image, eliminating the requirement for per-instance optimization or densely captured images. This empowers our method to generalize effectively in diverse scenarios while delivering high-quality reconstructions.
@@
The key insight behind our method is to reconstruct Gaussian properties from the diffusion latent space in a generalizable end-to-end architecture, integrating a 2D generative diffusion model as appearance prior and a human parametric model as structure prior. Specifically, facing this under-constraint reconstruction problem from single-view input with extensive invisible regions, we first leverage a 2D multi-view diffusion model (denoted as novel-view synthesizer) to hallucinate the unseen parts of clothed humans. Then a generalizable latent reconstruction Transformer is proposed to enable interaction among the generated diffusion latent embeddings and the geometric information of the structured human model, enhancing the quality of 3DGS reconstruction. During interactions, to mitigate the limitations of inaccurate human priors like the SMPL model, we devise a projection strategy to strike a balance between robustness and flexibility, projecting 3D tokens into the 2D space and conducting searches within adjacent windows using projection-aware attention. Another specific challenge of human reconstruction is to capture fine detail in visually sensitive areas, such as the face and hand. To address this problem, we exploit the semantic cues from the structure priors and propose semantics-guided objectives to further promote the fine details reconstruction quality. By synergistically amalgamating these crucial components into a unified framework, our method achieves state-of-the-art performance in striking the right balance between quality and efficiency.
@@
Figure 1: Our method achieves state-of-the-art rendering quality while maintaining the fastest runtime. (a) Qualitative results: LGM [29] and GTA [30] are generalizable but in lower quality, TeCH [21] exhibits issues with multi-face rendering and is time-consuming. In contrast, our method achieves higher fidelity in a much shorter time. (b) Performance and runtime comparison: metrics are evaluated on the challenging Twindom dataset.
@@ 2 Related Work
Single-Image Human Reconstruction. Single-image human reconstruction methods fall into explicit and implicit approaches. Explicit methods use parametric body models [5, 6] to estimate minimally clothed human meshes [1, 2, 31, 32, 33, 34, 35, 36, 37], adding clothing via 3D offsets [3, 4, 38, 39, 40, 41] or garment templates [42, 43, 44]. However, these methods are constrained by topology, particularly with loose clothing. Implicit methods [7, 8, 9, 45, 22, 10, 11, 27, 46], utilizing representations like occupancy, signed distance fields (SDF) and Neural Radiance Fields (NeRF), offer flexibility with topology, allowing for accurate depiction of 3D clothed humans. Recent methods [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 30, 57, 58, 59, 60, 61] combine implicit representation with explicit human priors for better robustness. Among these, GTA [30] uses Transformers with fixed embeddings for translating image features into 3D tri-plane features. TeCH [62] employs diffusion-based models for visualizing unseen areas but needs extensive optimization and accurate SMPL-X models. Another approach, HumanSGD [22], uses diffusion models for texture inpainting but struggles with mesh inaccuracies from [8]. NeRF-based methods like SHERF [10] generate human NeRF from single images using hierarchical features, while ELICIT [11] leverages CLIP [63] for contextual understanding. These NeRF-based methods produce high-quality images but often struggle with detailed 3D mesh generation and require extensive optimization time.
@@
Human Gaussian Splatting. While implicit representations like SDF or NeRF struggle with balance optimization efficiency and rendering quality, the high efficiency of 3DGS [14] has advanced 3D human creation. Recent 3DGS methods have targeted multi-view videos, monocular video sequences, and sparse-view input. Multi-view video methods like D3GA [15], HuGS [16], and 3DGS-Avatar [17] exploit dense spatial and temporal information for detailed models. Additionally, HiFi4G [64] combine 3DGS with a dual-graph mechanism to preserve spatial-temporal consistency, ASH [65] utilizes mesh UV parameterization for real-time rendering, and Animatable Gaussians [66] improve garment dynamics via pose projection mechanism and 2D CNNs. The monocular video only provides sufficient observation as humans move, so the monocular human Gaussian Splatting methods [18, 19, 20, 67, 68, 69, 70, 71, 72, 73] often resort to using LBS for mapping to a canonical space and introducing additional regularization terms. For sparse-view images, GPS-Gaussian [28] achieves high rendering speeds without per-subject optimization by using a feed-forward way with efficient 2D CNNs encoding from diverse 3D human scan data. In comparison, our work stands out as the first one-shot generalizable human GS method, using only single image input and achieving high-quality reconstruction without any optimization and fine-tuning.
@@
Generalizable Large Reconstrucion Model. Large reconstruction model (LRM) [74, 75] presents that with sufficient parameters and training datasets, a deterministic feed-forward Transformer [76] is capable of learning a triplane feature for volumetric rendering from a single-view image. Targeting general 3D object, TriplaneGaussian [77] combines 3DGS with triplane as a hybrid 3D representation for fast rendering, where point clouds are first extracted from an image and then projected on triplane features to decode 3DGS attributes. Most recent methods further leverage 2D diffusion models that can generate multi-view images simultaneously [78, 79, 80, 81] to provide plausible and abundant inputs for the following reconstruction, so they focus on the sparse multi-view reconstruction task. Instant3D [81] follows the technique of LRM and produces triplane features from 4 posed images by a Transformer encoder. CRM [82], MeshLRM [83] and InstantMesh [84] replaces triplane NeRF by FlexiCubes [85] to extract meshes directly. LGM [29], GRM [86] and GS-LRM [87] predict 3DGS attributes from each input pixel [88, 89] and combine the Gaussians from multiple viewpoints as the final 3D output. For human task, CharacterGen [90] follows Instant3D [81] to adopt a two-stage approach: first, generating a 2D multi-view canonical images, then using multi-view SDF reconstruction for cartoon digital characters, without direct interaction between stages. A more compact and elegant method HumanLRM [27] replaces training data of LRM [74] from general objects with human data to predict triplane NeRF, showcasing the capacities of the large model. While it relies heavily on the dataset, often leading to issues like missing limbs due to a lack of prior knowledge. In contrast, our 3DGS-based method leverages the latent diffusion and reconstruction model in an end-to-end framework and introduces human priors by addressing their inaccuracy with specific projection strategies, achieving better generalization and stability with even less data.
@@ 3 Method
Figure 2: Overview of HumanSplat. (a) Multi-view latent features are first generated by a fine-tuned multi-view diffusion model (Novel View Synthesizer in Sec. 3.3). (b) Then, the Latent Reconstruction Transformer (Sec. 3.4) interacts global latent features (Sec. 3.4.1) and human geometric prior (Sec. 3.4.2). (c) Finally, the semantic-guided objectives (Sec. 3.5) are proposed to reconstruct the final human 3DGS.
@@ 3.1 Preliminary
SMPL [5] is a widely-used parametric human body model that is created by skinning and blend shapes, and learned from thousands of 3D body scans. The SMPL model [5] utilizes shape parameters $\boldsymbol{\beta}\in\mathbb{R}^{10}$ and pose parameters $\boldsymbol{\theta}\in\mathbb{R}^{24\times 3}$ to parameterize the deformation of the human body mesh $\mathcal{M}(\boldsymbol{\beta},\boldsymbol{\theta}):\boldsymbol{\beta}\times\boldsymbol{\theta}\mapsto\mathbb{R}^{6890\times 3}$.
3D Gaussians Splatting [14] is a recently proposed 3D representation that formulates 3D content by a set of $\mathbf{N_{p}}$ colored Gaussians $\mathcal{G}=\{\mathbf{g}_{i}\}_{i=1}^{\mathbf{N_{p}}}$. Each Gaussian $\mathbf{g}_{i}=\exp\left(-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu}_{i})^{\top}\boldsymbol{\Sigma}_{i}^{-1}(\mathbf{x}-\boldsymbol{\mu}_{i})\right)$ has opacity $\boldsymbol{\sigma}_{i}\in\mathbb{R}$ and color $\mathbf{c}_{i}\in\mathbb{R}^{3}$ attributes, where $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$ is its location and $\boldsymbol{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$ implies its scaling $\mathbf{s_{i}}\in\mathbb{R}^{3}$ and orientation quaternion $\mathbf{q_{i}}\in\mathbb{R}^{4}$ in 3D space. These attributes constitute $\mathbf{G}\in\mathbb{R}^{\mathbf{N_{p}}\times 14}$ can represent a 3D object, and by projecting to the image plane with opacity-based color composition in the depth order, it can be differentiably rendered into 2D images in real-time.
@@ 3.2 Overview
Given a single input image of a human body $\mathbf{I_{0}}\in\mathbb{R}^{H\times W\times 3}$, our task is to reconstruct a 3D representation, i.e., 3DGS in this work, from which novel-view images can be rendered. As illustrated in Fig. 2, our method leverages a 2D diffusion model and a novel latent reconstruction Transformer that adeptly integrates 2D appearance priors, human geometric priors, and semantic cues within a universal framework. First, we use SMPL estimator [91, 92] to predict human structure priors $\mathcal{M}$, and CLIP [63] to generate the image embedding of the input image $\mathbf{c}$. A latent temporal-spatial diffusion model [93], referred to as the novel-view synthesizer (Sec. 3.3), is fine-tuned and produces multi-view latent features $\left\{\mathbf{F}_{i}\in\mathbb{R}^{h\times w\times c}\right\}_{i=1}^{\mathbf{N}}$, where $\mathbf{N}$ is the number of views, and $h$, $w$ and $c$ are the height, width and channel of the latent features. Then, we utilize a novel latent reconstruction Transformer (Sec. 3.4) that adeptly integrates human geometric prior and latent features within a universal framework, which predicts Gaussian attributes $\mathbf{\mathbf{G}}\coloneqq\left\{(\boldsymbol{\mu}_{i},\boldsymbol{q}_{i},\boldsymbol{s}_{i},\boldsymbol{c}_{i},\boldsymbol{\sigma}_{i})|i=1,...,\mathbf{N_{p}}\right\}$ and rendered into new images, where $\mathbf{N_{p}}$ represents the number of Gaussian points. Furthermore, to achieve high-fidelity texture modeling and better constrain the estimated multiple views, we designed a hierarchical loss that incorporates human semantic priors (See 3.5). Note that during training, the network is trained using 3D scan data to ensure accurate supervision from various viewpoints, where the registered SMPL is fitted via the multi-view version of SMPLify [92]. During inference, the model employs PIXIE [91] to predict human structure priors $\mathcal{M}$ and synthesize novel views from only a single image input based on the trained model, without any fine-tuning and optimization.